# An explicit construction of Kaleidocycles by elliptic theta functions

### v3

@@ 5.2. Motion of vertex positions
(5.96) $\begin{split}&Q_{n}Q_{n}^{*}=4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
@@
Furthermore, if the left-hand-side of the equation (5.136) does not depend on $n$, the following holds:
@@
The proofs of (i) and (ii) follow from the same calculation as in Section 5 of [18]. We prove (iii). From the deformation equation (5.103) and the discrete Frenet-Serret formula (2.5), we see that
@@
holds, which vanishes because of (5.136). Therefore $\dot{\lambda}=0$.
@@
Proposition 5.9 (iii) does not hold in general without the assumption on (5.136). Indeed, there is a motion of a Kaleidocycle satisfying (5.103) such that the left-hand side of (5.136) depends both on $n$ and $t$.
@@
(5.109) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0,\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}=0.\end{split}$
@@ 6. Existence of Kaleidocycles
The constant torsion-angle and unit segment-length curves given in Theorem 5.1 are not necessarily closed. In this section, we derive a necessary and sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle. Furthermore, we prove that there exist parameters that satisfy the closure condition when $k\geq 6$. The overall approach is as follows: For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle has to have period $k$. From this, we see that $m:=kv/y$ is an integer (§6.1). Then, rewriting the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ assuming $m$ is an integer, we obtain two equations for v,r,yv,r,y6.2). By rearranging one of the equations, we verify that $r=r(y)$ is uniquely determined when $y$ is in a certain range (§6.3). Substituting this $r(y)$ into the other equation, we find that if $m\geq 3$, then there exists $y$ satisfying it (§6.4). By considering two transformations of $m$ and $n$ that preserve the shape of a curve, we see that it is sufficient to consider $0<m\leq k/2$. Finally, we show that under this range of $m$ and $k\geq 6$, there always exists a set of parameters with which the curve Theorem 5.1 is closed, and hence, gives a $1$-dimensional periodic motion of a Kaleidocycle (§6.5).
@@ 6.1. Period of the curvature angle
We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$. In this case, the absolute value of the curvature angles also has a period $k$. Hence, we obtain the necessary condition for the curve to be closed that
@@
holds for all $n$ and $t$. Note that there is no Kaleidocycle with period $1$, hence $k\geq 2$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
@@
where we note $\tilde{\mu}_{n}(t)=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}$. Since $\tilde{\mu}_{n}(t)$ can be regarded as a linear function of $t$ when $v,y$ and $n$ are fixed, $\tilde{\mu}_{n}(t)$ attains all reals as $t$ varies. Thus, we only need to check that (6.2) holds for $n=0$. Since $\mathrm{cn}^{2}$ has a real period $2\tilde{K}$, $m=kv/y$ must be an integer. Hereafter, let $v$ satisfy
@@ 6.2. Closure conditions
Here we obtain the necessary and sufficient conditions satisfied by $v,r,y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
@@
We rewrite the left-hand side further with (A.12) under the condition that $m$ is an integer. Then we obtain
@@
On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v,r,y$ satisfy $kv=my$, (6.6) and (6.10). Thus, they give necessary and sufficient condition for the curve given by Theorem 5.1 to be a closed curve for every $t\in\mathbb{R}$.
6.3. An explicit solution to the condition (6.10)
@@
where $\xi=\vartheta_{2}\left(0\right)^{2}/\vartheta_{3}\left(0\right)^{2}$ is an elliptic modulus and $K=\pi\vartheta_{3}\left(0\right)^{2}/2$ is a complete elliptic integral of the first kind. We define the following function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
@@
When $1<\varphi\left(m/k,y\right)<1$ is satisfied, $r=r(y)$ is uniquely determined because the inverse function of $\mathrm{cn}$ exists.
6.3.1. Existence of a inverse function
@@
(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<1/2).$
@@
$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<1/2),$
@@
(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<1/2).$
@@ 6.4. Existence of a solution to the condition $\Lambda=0$
By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$. Note that $\Lambda+1$ can also be written as
@@
using a real-valued function $\Psi=\Psi(v,r,y)$. When
@@
is greater than $2\pi$, then there exists a solution to $\Lambda(v,r(y),y)=0$ in the range $0<y<y_{0}(m/k)$. Here, we compute the above quantities and show that such a $y$ exists if $m\geq 3$. From a representation by an infinite product of elliptic theta functions,
@@
Furthermore, from
@@
we obtain
@@
Therefore, when $m\geq 3$, there is always a solution in $\Lambda(v,r(y),y)=0$.
@@ 6.5. Transformations of a parameter and a variable
In the following, we will construct the self-bijections and prove that the Kaleidocycles obtained by the self-bijections are essentially identical to the Kaleidocycles before the transformation. A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarised in Table 1.
@@ 6.5.1. Transformation $m\rightarrow m+k$, $n\rightarrow k-n$
That is, the new Kaleidocycle is obtained by shifting the time parameter, reversing the orientation of the central curve, and flipping every other hinge, which is equivalently encoded by the torsion angle update $\lambda\mapsto\pi-\lambda$.
@@ 6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
Summarising the results of this section, we have the following theorem.
@@
For a fixed $k\geq 6$, let $0<m\leq k/2$. Then there exist $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $y>0$ satisfying $kv=my$, (6.6) and (6.10). Then, the curve defined by Theorem 5.1 is closed for every $t\in\mathbb{R}$ and forms a $k$-Kaleidocycle with the torsion angle given by (5.2).
@@
Table 1 summarises the relationships among the three deformation equations.
@@
$\downarrow$ Specialise the deformation equation by choosing
@@ Appendix A Functional identities of elliptic theta functions
In the appendix, we collect some properties of elliptic theta functions which we rely heavily on our calculations.

### v4

@@ 5.2. Motion of vertex positions
(5.97) $\begin{split}Q_{n}Q_{n}^{*}=&4\beta_{3}\left(\beta_{1}-\beta_{3}\right)F_{n}^{4}+\left\{1+2\left(\beta_{1}-2\beta_{3}\right)\left(\beta_{4}+\beta_{4}^{*}\right)\right\}F_{n}^{2}|G_{n}|^{2}\\ &-\left(\beta_{4}+\beta_{4}^{*}\right)^{2}|G_{n}|^{4}.\end{split}$
@@
Furthermore, assume that
(5.109) $\tan\frac{\kappa_{n+1}}{2}\tan\frac{\kappa_{n-1}}{2}-\tan^{2}\frac{\kappa_{n}}{2}+\tan\frac{\kappa_{n+1}}{2}\tan^{2}\frac{\kappa_{n}}{2}\tan\frac{\kappa_{n-1}}{2}$
is independent of $n$ (see 5.14). Then,
@@
The proofs of (1) and (2) follow from the same calculation as in Section 5 of [18]. We prove (3). From the deformation equation (5.104) and the discrete Frenet-Serret formula (2.5), we see that
@@
holds. Put $a_{n}=\tan\frac{\kappa_{n}}{2}$. Using
$\sin\kappa_{n}=\frac{2a_{n}}{1+a_{n}^{2}},\qquad\tan\kappa_{n}=\frac{2a_{n}}{1-a_{n}^{2}},$
the bracket in (5.110) becomes
$\frac{1}{2}\Bigl(a_{n+2}a_{n}-a_{n+1}^{2}+a_{n+2}a_{n+1}^{2}a_{n}-a_{n+1}a_{n-1}+a_{n}^{2}-a_{n+1}a_{n}^{2}a_{n-1}\Bigr).$
Hence it is equal to one half of the difference between (5.109) at indices $n+1$ and $n$. Therefore it vanishes if the quantity in (5.109) is independent of $n$, implying $\dot{\lambda}=0$.
@@
Proposition 5.9 (3) does not hold in general without the assumption on (5.109). Indeed, there is a motion of a Kaleidocycle satisfying (5.104) such that (5.109) depends both on $n$ and $t$.
@@
(5.111) $\begin{split}\Gamma_{+}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)+\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{+}\left(v,r,y\right)=&-i\Gamma_{+}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)\\ &-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}},\\ \rm or\\ \Gamma_{-}\left(r,y\right)=&\frac{\vartheta_{1}^{{}^{\prime}}(2r)\vartheta_{3}(0)-\vartheta_{3}\left(2r\right)\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}\left(2r\right)\vartheta_{3}\left(0\right)},\\ C_{-}\left(v,r,y\right)=&-i\Gamma_{-}\left(\beta_{2}\left(2\beta_{3}-\beta_{1}\right)-\frac{\beta_{4}+\beta_{4}^{*}}{\beta_{2}}\right)\\ &-\beta_{6}-\left(\beta_{5}+\beta_{5}^{*}\right)\left(2\beta_{3}-\beta_{1}\right)+\frac{\beta_{5}-\beta_{5}^{*}}{\beta_{2}}.\end{split}$
@@ 6. Existence of Kaleidocycles
The curves given in Theorem 5.1 are not necessarily closed. In this section, we derive necessary and sufficient conditions on the parameters $v$, $r$, and $y$ for the curve $\gamma$ to be closed, and hence to define a Kaleidocycle. We also prove that such parameters exist when $k\geq 6$.
The overall strategy is as follows. First, we derive a necessary condition for closedness. For the curve $\gamma$ to be closed, the absolute value of the signed curvature angle must be periodic with period $k$. This implies that $m:=kv/y\in\mathbb{Z}\setminus(k\mathbb{Z})$6.1).
Assuming $m\in\mathbb{Z}\setminus(k\mathbb{Z})$, we then rewrite the closedness condition $\gamma_{n+k}(t)=\gamma_{n}(t)$ as two equations in the parameters $v$, $r$, and yy6.2). One of these equations determines $r=r(y)$ uniquely when $y$ lies in a certain range (§6.3). Substituting this expression into the remaining equation, we show that a solution $y$ exists whenever $3\leq m\leq k/2$6.4).
Finally, by considering two transformations of $m$ and $n$ that preserve the shape of the curve, we may assume without loss of generality that $3\leq m\leq k/2$. Combining these results, we conclude that for $k\geq 6$, there always exists a choice of parameters for which the curve in Theorem 5.1 is closed, thereby establishing the existence of a Kaleidocycle (§6.5).
@@ 6.1. Period of the curvature angle
We say a curve $\gamma$ is closed with a period $k\geq 2$ when $\gamma_{n+k}(t)=\gamma_{n}(t)$ for all $n$ and $t$. In this case, the absolute value of the curvature angles also has a period $k$. Hence, we obtain the necessary condition for the curve to be closed that
@@
holds for all $n$ and $t$. Rewriting this condition using Jacobi $\mathrm{cn}$-function, we obtain
@@
where we note $\tilde{\mu}_{n}(t)=\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}$. Since $\tilde{\mu}_{n}(t)$ can be regarded as a linear function of $t$ when $v,y$ and $n$ are fixed, $\tilde{\mu}_{n}(t)$ attains all reals as $t$ varies. Thus, we only need to check that (6.2) holds for $n=0$. Since $\mathrm{cn}^{2}$ has the real period $2\tilde{K}$, it follows that $m=kv/y$ must belong to $\mathbb{Z}\setminus(k\mathbb{Z})$. Hereafter, let $v$ satisfy
@@ 6.2. Closure conditions
Here we obtain the necessary and sufficient conditions satisfied by $v$,$r$, and $y$ for the curve to be closed. We first rewrite $\gamma_{n+k}(t)=\gamma_{n}(t)$ in terms of $\tau$ functions. We have
@@
We further rewrite the left-hand side using $m$ and (A.12). Then we obtain
@@
On the other hand, the closure condition (6.4) holds for any $n\in\mathbb{Z}$ and $t\in\mathbb{R}$ when the parameters $v$, $r$, and $y$ satisfy $kv=my$, (6.6) and (6.10). Thus, they give necessary and sufficient condition for the curve given by Theorem 5.1 to be a closed curve for every $t\in\mathbb{R}$.
6.3. An explicit solution to the condition $\frac{\partial}{\partial r}\log(\Lambda+1)=0$
@@
where $\xi=\vartheta_{2}\left(0\right)^{2}/\vartheta_{3}\left(0\right)^{2}$ is an elliptic modulus and $K=\pi\vartheta_{3}\left(0\right)^{2}/2$ is a complete elliptic integral of the first kind. We define the following continuous real-valued function $\varphi$ for the ratio of $m$ and $k$ and for $y>0$.
@@
When $1<\varphi\left(m/k,y\right)<1$ is satisfied, (6.13) determines a unique $r=r(y)$ in $(0,1/2)$ because on the interval $0<r<1/2$, the function $r\mapsto\mathrm{cn}(4Kr,\xi)$ is strictly decreasing from $1$ to 1-1.
6.3.1. Existence of an inverse function
@@
(6.17) $\lim_{y\to+0}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to+0}\frac{\vartheta_{2}\left(0;\frac{i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}{\vartheta_{4}\left(0;\frac{i}{y}\right)\vartheta_{2}^{{}^{\prime}}\left(\frac{m}{k};\frac{i}{y}\right)}=0\quad(0<\frac{m}{k}<\frac{1}{2}).$
@@
$ip^{\left(\frac{m}{k}\right)^{2}}\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y;iy\right)$ $=-2\pi p^{\left(1-\frac{m}{k}\right)^{2}}+\mathcal{O}(p^{\left(1+\frac{m}{k}\right)^{2}})\quad(0<\frac{m}{k}<\frac{1}{2}),$
@@
(6.19) $\lim_{y\to\infty}\varphi\left(\frac{m}{k},y\right)=\lim_{y\to\infty}\frac{\vartheta_{4}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{2}\left(i\frac{m}{k}y\right)+\vartheta_{2}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}{\vartheta_{2}\left(0\right)\left(2\pi i\frac{m}{k}\vartheta_{4}\left(i\frac{m}{k}y\right)+\vartheta_{4}^{{}^{\prime}}\left(i\frac{m}{k}y\right)\right)}=-\infty\quad(0<\frac{m}{k}<\frac{1}{2}).$
@@ 6.4. Existence of a solution to the condition $\Lambda=0$
By substituting $r=r(y)$ obtained in the previous section into $\Lambda(v,r,y)$, we prove that when $3\leq m\leq k/2$, $\Lambda(v,r(y),y)=0$ has a solution in $0<y<y_{0}$. Note that $\Lambda+1$ can also be written as
@@
using a continuous real-valued function $\Psi=\Psi(v,r,y)$. When
@@
is greater than $2\pi$, then there exists a solution to $\Lambda(v,r(y),y)=0$ in the range $0<y<y_{0}(m/k)$. Here, we compute the above quantities and show that such a $y$ exists if $3\leq m\leq k/2$. From a representation by an infinite product of elliptic theta functions,
@@
From the explicit expression of $r$ in (6.20), it follows that $1/4\leq r<1/2$ when $0<m\leq k/2$, which yields the following limit.
@@
This calculation remains valid even when $r$ is determined by (6.21), that is, in the case where $1/2<r\leq-1/4$. Therefore, we obtain
@@
Therefore, when $3\leq m\leq k/2$, there is always a solution in $\Lambda(v,r(y),y)=0$.
@@ 6.5. Transformations of a parameter and a variable
In the following, we will construct the self-bijections and prove that the Kaleidocycles obtained by the self-bijections are essentially identical to the Kaleidocycles before the transformation. A combination of the two bijection exchanges a solution of the semi-discrete potential mKdV equation with one of the semi-discrete sine-Gordon equation, and the results are summarized in Table 1.
@@ 6.5.1. Transformation $m\rightarrow m+k$, $n\rightarrow k-n$
That is, the new Kaleidocycle is obtained by shifting the time parameter, reversing the orientation of the central curve, and flipping every other hinge, which is equivalently encoded by the torsion angle update $\lambda\mapsto\pi-\lambda$. The argument remains valid when $r$ is calculated by (6.20).
@@ 6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
Summarizing the results of this section, we have the following theorem.
@@
For $k\geq 6$ and an integer satisfying $3\leq m\leq\frac{k}{2}$, there exist $y>0$, $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ such that $kv=my$ and (6.6) and (6.10) hold. Hence the curve given in Theorem 5.1 is closed for every $t\in\mathbb{R}$ and defines a $k$-Kaleidocycle whose torsion angle is given by (5.2).
@@
Table 1 summarizes the relationships among the three deformation equations.
@@
$\downarrow$ Specialize the deformation equation by choosing
@@ Appendix A Functional identities of elliptic theta functions
In this appendix, we collect several properties of elliptic theta functions on which our calculations rely heavily.
@@ A.5. Bilinear identities connecting theta functions with $y$ and $y/2$
Conflicts of Interest
The authors declare no conflicts of interest.
Data Availability Statement
The computer program used in the experiments is available in the associated GitHub repository https://github.com/shizuo-kaji/Kaleidocycle/.